# StructuredField: Unifying Structured Geometry and Radiance Field

### v1

@@ Abstract
Recent point-based differentiable rendering techniques have achieved significant success in high-fidelity reconstruction and fast rendering. However, due to the unstructured nature of point-based representations, they are difficult to apply to modern graphics pipelines designed for structured meshes, as well as to a variety of simulation and editing algorithms that work well with structured mesh representations. To this end, we propose StructuredField, a novel representation that achieves both a structured geometric representation of the reconstructed object and a high-fidelity rendering reconstruction of the object. We employ structured tetrahedral meshes to represent the reconstructed object. We reparameterize the geometric parameters of the tetrahedral mesh into the geometric shape parameters of a 3D Gaussians, thereby achieving differentiable high-fidelity rendering of the tetrahedral mesh. We propose a novel inversion-free homeomorphism to constrain the optimization of the tetrahedral mesh, which strictly guarantees that the tetrahedral mesh is remains both inversion-free and self-intersection-free during the optimization process and the final result. Based on our proposed StructuredField, we achieve high-quality structured meshes and high-fidelity reconstruction. We also demonstrate the applicability of our representation to various applications such as physical simulation and deformation. Project page: https://structuredfield.github.io
1 Introduction
@@
Figure 1: StructuredField represents and reconstructs scene using a structured tetrahedral mesh. This novel structured 3D representation enables a variety of applications, including physical simulations, deformations, and more.
@@
Photorealistic and real-time rendering of 3D scenes is a central pursuit in computer graphics, both in academic research and practical applications. Traditional 3D representations, such as meshes [9, 34, 19] and point clouds [47, 64], facilitate real-time rendering through rasterization techniques that align with modern GPU rendering architectures. Nevertheless, these methods frequently produce low-quality rendering, exhibiting missing geometric details and blurry artifacts. In contrast, emerging differentiable volumetric rendering methods [5, 41, 11], including neural radiance fields [39], demonstrate the capability to reconstruct 3D scenes in an end-to-end manner using multi-view images, thus achieving high fidelity and maintaining intricate details. However, the reliance on ray tracing-based rendering pipelines and the necessity for extensive sampling points pose challenges to rapid rendering in NeRF variants.
@@
Recently, advancements in point-based differentiable rendering techniques [30, 18, 23, 22] have demonstrated notable success in enhancing both rendering speed and high-fidelity reconstruction by leveraging smooth primitives alongside rasterization-based pipeline. Although point-based rendering techniques offer numerous advantages, they inherently represent scenes as unstructured point clouds. In contrast, contemporary graphics pipelines are designed for structured representations, such as polygonal meshes, which are increasingly favored in a variety of applications, including animation [25, 49], physical simulation [3, 4, 35] and editing [65, 68]. The lack of a representation for the structure of objects within point-based rendering methods significantly constrains their broader applications.
@@
In this study, our aim is to bridge the gap between recent high-fidelity differentiable point-based representations and traditional graphics pipelines that are tailored for structured meshes. To achieve this, we introduce a structured representation that facilitates both high-fidelity rendering and the recovery of near-surface structures. Specifically, we use tetrahedral mesh as the fundamental representation. Compared to triangular mesh, tetrahedral mesh is better suited for volumetric data and has been widely utilized in volume rendering [36, 61, 58]. We leverage point-based 3D representations as differentiable renderers and employ differentiable mappings to reparameterize primitive parameters with the attributes of tetrahedral mesh. Thereby, we enable the optimization of tetrahedral mesh parameters in a differentiable manner and achieve high-fidelity reconstruction.
However, we find that directly optimizing tetrahedral mesh without constraints leads to anomalous topological structures. To mitigate this challenge, we analyze two main causes of mesh anomalies: self intersections and element inversions. We employ homeomorphic mappings [12, 8] to prevent self intersections and introduce permutations to ensure that the Jacobian determinant remains positive, thereby preventing element inversions in mesh. By utilizing this orientation-preserving homeomorphism as a constraint within the feasible solution space, we can prevent the occurrence of self intersections or inversions in the tetrahedral mesh during the optimization process.
@@
  • We introduce a novel 3D structured representation that employs reparameterization to make tetrahedral mesh differentiable, facilitating high-fidelity scene reconstruction.
@@
  • We propose a novel invertible network architecture specifically designed to represent orientation-preserving homeomorphisms, which ensures that the tetrahedral mesh remains inversion-free and self-intersection-free during optimization.
  • StructuredField can achieve high quality geometry and rendering reconstruction while supporting physics simulation and deformation without any modifications.
2 Related Work
Figure 2: Overview of StructuredField. Given multi-view images as input, we reconstruct the 3D scene using a structured tetrahedral mesh. We use an orientation-preserving homeomorphism to maintain the inversion-free structured geometry of the tetrahedral mesh during optimization. Additionally, parameters of tetrahedral mesh are reparameterized into the corresponding 3D Gaussian parameters, enabling differentiable optimization. Finally, we use differentiable rasterization to render images from the given viewpoints. Based on our structured mesh representation, we can further apply the model to physical simulations, deformation, and other applications.
2.1 3D Reconstruction
@@ 2 Related Work
Reconstructing 3D scenes from multi-view images is a longstanding problem in both computer graphics and computer vision. Traditional 3D reconstruction techniques include Structure-from-Motion (SfM) pipelines [2, 13] to estimate camera poses and obtain sparse point clouds, followed by surface reconstruction through dense multi-view stereo [16, 28, 15]. These methods rely on hand-crafted features to acquire fine textures and geometry, and struggle to reconstruct view-dependent colors. Significant advancements have been achieved in NVS, particularly since the introduction of Neural Radiance Fields (NeRF) [39]. The original NeRF represents the scene as an MLP, which maps positional encodings of spatial locations and directions to attributes including color and density, and utilizes volume rendering process to achieve realistic rendering. Various works have enhanced the performance of NeRF [41, 7, 5] or extended them to large scenes [6, 50]. More recently, 3DGS [30] optimizes anisotropic 3D Gaussian primitives, demonstrating real-time photorealistic reconstruction results. This method has been quickly extended to multiple domains [33, 29, 51, 60]. Despite these successes, point-based representations are unstructured, limiting their further applications. In this paper, we demonstrate detailed reconstruction while maintaining a structured tetrahedral mesh, and showcase its subsequent applications.
2.2 Mesh-based Representation
@@
Explicit representations have served as a cornerstone within 3D modeling and computer graphics for decades [59]. Conventional geometric representations like point clouds, voxels and polygonal meshes have been extensively revisited in the context of 3D deep learning. Polygonal meshes are particularly attractive due to their structured geometry and efficient rendering properties. Recently, differentiable rendering methods [34, 43] leveraging mesh representations have enabled the production of high-quality renderings. However, the optimization processes inherent to mesh representations are often hampered by rigid topological constraints, resulting in limited flexibility and reduced capacity to accurately depict realistic appearances. Our method integrates inversion-free mapping to effectively regulate and maintain structural integrity during the mesh optimization process while ensuring high quality reconstruction, addressing these limitations.
@@
On the other hand, NeRF [38] has attracted considerable attention due to their ability to deliver high-fidelity renderings. Substantial research has focused on integrating implicit methods with mesh structures to extend and refine NeRF [54]. For instance, some approaches [57, 53, 52] utilize meshes to constrain the sampling regions within an adaptively learned narrow ban of two explicit meshes. On the other hand, Tetra-NeRF [31, 46] employs tetrahedral mesh as feature grids to accelerate the training process. Additionally, leveraging the fast rendering capabilities of meshes, Baked SDF and its variants [63, 45] have baked radiance fields into meshes to facilitate real-time rendering. Mesh representations have also been pivotal in the deformation of radiance fields. NeRFShop and Cage-NeRF [42, 24] using meshes as cages to drive the deformation of radiance fields.
@@
Recently, point-based representations [30, 23, 22] have emerged, utilizing smooth primitives to represent radiance fields and employing rasterization techniques to accelerate rendering. In the subsequent works, Mani-gs and GaMes [17] employ high-quality triangle meshes or triangle soups as proxies, binding primitives to the triangle mesh to drive the deformation of 3D Gaussians. Additionally, VR-GS [26] and D3GA [67] adopt tetrahedral mesh as proxies to facilitate the deformation of objects. However, these methods directly bind the mesh to Gaussians without optimizing the positions of the mesh vertices. As a result, any imperfections in the initial mesh can significantly degrade the final reconstruction quality. In contrast, our method reparametrizes Gaussians with tetrahedral mesh vertices and optimizes these vertices, thereby allowing for high-qualityreconstruction.
2.3 Inversion-free Mesh Optimization
@@
Maintaining mesh quality during the optimization process remains a challenging problem. Element inversion is a major factor causing mesh quality issues [14]. Traditional methods [48, 44] impose constraints on the Jacobian of the mapping during optimization to ensure the Jacobian remains positive, thereby enforcing that the mapping is orientation-preserving. In addition to being inversion-free, intersection-free boundaries are also necessary during the optimization. This is typically achieved by constraining the mapping to be bijective. Traditional methods [27, 40] introduce scaffold meshes to convert the globally overlap-free constraint into a locally flip-free condition. Recently, invertible networks [8, 12] offer another solution, leading to their widespread adoption in recent 3D deformation tasks. For instance, NDR [10] and Cadex [32] employ invertible networks to model the motion of objects, thereby achieving dynamic object reconstruction. NFGP [62] utilizes invertible networks to perform geometric processing on implicit surfaces. Compared to these methods, we propose a novel invertible network that implicitly and strictly constrains the mapping to be inversion-free and intersection-free, and apply it during the optimization process to guarantee the mesh quality.
3 Method
@@
Given multi-view posed images of a scene, our primary objective is to reconstruct it with high fidelity, capturing not only its geometric details but also the structures near its surface. This enables seamless integration of the reconstructed 3D scene into existing computer graphics pipelines, such as those used for rendering, animation, physical simulation, and deformation. To this end, we first introduce a novel tetrahedral mesh-based 3D representation that is structurally organized and facilitates differentiable optimization (Sec. 3.1). The mesh is optimized using orientation-preserving homeomorphism to ensure high-quality structured geometry (Sec. 3.2). An overview of our representation and reconstruction pipeline is shown in Fig. 2.
@@
Given a tetrahedral mesh $\mathcal{M}=(\mathcal{V},\mathcal{T})$, where $\mathcal{V}=\{v_{i}\}_{i=1}^{N}$ represents the set of vertices, $\mathcal{T}=\{T_{k}\}_{k=1}^{K}$ is the set of tetrahedron. Each tetrahedron $T_{k}\in\mathcal{T}$ is defined by four vertices $\{v_{k_{1}},v_{k_{2}},v_{k_{3}},v_{k_{4}}\}$. Tetrahedral mesh is generally challenging to render in a differentiable manner due to their discrete, non-smooth nature.
@@
To address this limitation, we build upon recent point-based differentiable rendering techniques, such as 3D Gaussian Splatting (3DGS) [30], 2D Gaussian Splatting (2DGS) [22], and convex splatting [18]. In our method, we establish a one-to-one correspondence between each tetrahedron and a primitive, using a differentiable reparameterization function $\mathbf{F}_{r}(\cdot)$ to map the parameters $T_{k}$ of the tetrahedron to the parameters $\Theta_{k}$ of the primitive:
$\Theta_{k}=\mathbf{F}_{r}\bigl{(}T_{k}\bigr{)}.$ (1)
@@
Gradients are propagated back to the parameters of tetrahedra through the differentiable reparameterization function, enabling the optimization of vertex positions and associated attributes in a fully differentiable framework. In the following, we demonstrate how the parameters of 3DGS can be reparameterized using the parameters of a tetrahedral mesh.
Figure 3: 2D illustration of reparameterizing Gaussians. In (a), (c), and (d), all vertex weights are equal, $w_{1}=w_{2}=w_{3}=1.0$. In (b), the bottom vertices $v_{2}$ and $v_{3}$ have higher weights, $w_{1}=0.1,w_{2}=w_{3}=1.0$.
3.1 Representation of StructuredField
@@
3DGS [30] represent a scene with a set of 3D Gaussians $\mathcal{G}=\{\boldsymbol{g}_{k}\}_{k=1}^{K}$, where each Gaussian $\boldsymbol{g}_{k}$ encodes the following attributes: mean $\boldsymbol{\mu}_{k}\in\mathbb{R}^{3}$, scales $\boldsymbol{s}_{k}\in\mathbb{R}^{3}$, rotation $\boldsymbol{r}_{k}\in\mathbb{R}^{4}$, color $\mathbf{c}_{k}\in\mathbb{R}^{3}$, and opacity $o_{k}\in\mathbb{R}$. In our representation, each tetrahedron $T_{k}$ contains four vertices with positions $\{\mathbf{v}_{k_{i}}\}_{i=1}^{4}$, spherical harmonic coefficients $\{\mathbf{SH}_{k_{i}}\}_{i=1}^{4}$, and weights $\{w_{k_{i}}\}_{i=1}^{4}$. We adopt a PCA-style [1] approach to convert these vertices and their attributes into a 3D Gaussian. Specifically, the mean and covariance matrix of each Gaussian are computed as:
$\boldsymbol{\mu}$ $=\frac{\sum_{i=1}^{4}w_{i}\,\mathbf{v}_{i}}{\sum_{i=1}^{4}w_{i}},$ (2)
$\boldsymbol{\Sigma}=$ $\frac{\sum_{i=1}^{4}w_{i}\bigl{(}\mathbf{v}_{i}-\boldsymbol{\mu}\bigr{)}\bigl{(}\mathbf{v}_{i}-\boldsymbol{\mu}\bigr{)}^{\top}}{\sum_{i=1}^{4}w_{i}}.$
This conversion effectively captures the weighted shape of the tetrahedron, ensuring that the resulting 3D Gaussian closely resembles the original tetrahedral shape, as shown in  Fig. 3.
We define the opacity attribute $o_{k}$ on the tetrahedron instead of on the vertices, so the opacity of its corresponding Gaussian is equivalent to the opacity of the tetrahedron. The color of the Gaussian is obtained by a weighted average of the colors computed from the spherical harmonic function of each vertex:
$\mathbf{c}=\frac{\sum_{i=1}^{4}w_{i}\mathbf{c}_{i}}{\sum_{i=1}^{4}w_{i}},\quad\mathbf{c}_{i}=\mathbf{SH}_{i}(\mathbf{d}),$ (3)
@@
After reparameterizing 3D gaussians with tetrahedral mesh, the scene can be rendered in a differentiable manner. However, freely optimizing the tetrahedral vertices leads to low quality tetrahedral mesh, as shown in  Fig. 4(a). In the following section, we introduce the orientation-preserving homeomorphism to address this issue.
Figure 4: Visualization of mesh optimization using different methods. We retain tetrahedra in the region where $y>0$ to better visualize the internal structure of the tetrahedral mesh. The experimental setup can be found in Sec. 4.1. The regularization loss used in the figure is $\mathcal{L}_{sv}$.
3.2 Inversion-free Structured Geometry
One of our goals is to optimize the tetrahedral mesh while maintaining high-quality structure. As shown in Fig. 4(a), optimizing a tetrahedral mesh without any constraints leads to poor mesh quality. The main cause of this issue is that the vertices can freely move during the optimization of the tetrahedral mesh. Consequently, the tetrahedral mesh suffers from self-intersections and element inversion problems, resulting in tetrahedra overlaps as shown in Fig. 5.
Figure 5: Two causes leading to mesh anomalies: self intersection and element inversion.
A naive solution is to introduce a loss function to constrain the mesh optimization process. Locally injective mapping [48] uses a barrier function to strictly prevent element inversion. However, the gradient explosion near the critical points of the barrier function makes it difficult to use directly in gradient descent-based optimization. Another approach is to apply a volume-preserving constraint $\mathcal{L}_{\text{volume}}$, or directly impose an $L_{1}$-norm constraint on the signed volume of the optimized tetrahedra $L_{\text{sv}}$, making tetrahedra with negative signed volume sparse:
$\mathcal{L}_{\text{volume}}$ $=\frac{1}{K}\sum_{k=1}^{K}\left\|s(T_{k})-s(T_{k}^{0})\right\|_{2},$ (4)
$\mathcal{L}_{\text{sv}}$ $=\frac{1}{K}\sum_{k=1}^{K}\max\left(-s(T_{k}),0\right),$ (5)
where $s(T_{k})$ is the signed volume of the tetrahedron $T_{k}$, $T^{0}_{k}$ is the initial tetrahedron. However, we found that these loss function-based constraints significantly impact the rendering quality and do not fully guarantee that the tetrahedral mesh remains reasonable structure as shown in Sec. 4.1.

### v2

@@ Abstract.
Recent point-based differentiable rendering techniques have achieved significant success in high-fidelity reconstruction and fast rendering. However, due to the unstructured nature of point-based representations, they are difficult to apply to modern graphics pipelines designed for structured meshes, as well as to a variety of simulation and editing algorithms that work well with structured mesh representations. To this end, we propose StructuredField, a novel representation that achieves both a structured geometric representation of the reconstructed object and high-fidelity rendering reconstruction. We employ structured tetrahedral meshes to represent the reconstructed object. We reparameterize the geometric attributes of these tetrahedra into the parameters of 3D Gaussian primitives, thereby enabling differentiable, high-fidelity rendering directly from the mesh. Furthermore, a hierarchical implicit subdivision strategy is utilized to ensure a conformal mesh structure while empowering the representation to capture multi-scale details. To maintain geometric integrity during optimization, we propose a novel inversion-free homeomorphism that constrains the tetrahedral mesh, guaranteeing it remains both inversion-free and self-intersection-free during the optimization process and in the final result. Based on our proposed StructuredField, we achieve high-quality structured meshes that are completely inversion-free and conformal, while also attaining reconstruction results comparable to those of 3DGS. We also demonstrate the applicability of our representation to various applications such as physical simulation, deformation, and level-of-detail. Code available at https://github.com/kevin2000A/StructuredField.
tetrahedral mesh, radiance fields, physics simulation, 3D reconstruction, differentiable rendering
submissionid: 1617journal: TOGccs: Computing methodologies Renderingccs: Computing methodologies Rasterizationccs: Computing methodologies Volumetric models
@@
Figure 1. StructuredField represents and reconstructs scene using a structured tetrahedral mesh. This novel structured 3D representation enables a variety of applications, including physical simulations, deformations, and more.
[…]
StructuredField represents and reconstruct scene using a structured tetrahedral mesh. This novel structured 3D representation enables a variety of applications, including physical simulations, deformations, level-of-detail and more.
1. Introduction
@@
Photorealistic and real-time rendering of 3D scenes is a central pursuit in computer graphics, both in academic research and practical applications. Traditional 3D representations, such as meshes (Botsch et al., 2005; Liu et al., 2019; Hoppe, 2023) and point clouds (Rusinkiewicz and Levoy, 2000; Yifan et al., 2019), facilitate real-time rendering through rasterization techniques that align with modern GPU rendering architectures. Nevertheless, these methods frequently produce low-quality rendering, exhibiting missing geometric details and blurry artifacts. In contrast, emerging differentiable volumetric rendering methods (Barron et al., 2021; Müller et al., 2022; Chen et al., 2022), including neural radiance fields (Mildenhall et al., 2020a), demonstrate the capability to reconstruct 3D scenes in an end-to-end manner using multi-view images, thus achieving high fidelity and maintaining intricate details. However, the reliance on ray tracing-based rendering pipelines and the necessity for extensive sampling points pose challenges to rapid rendering in NeRF variants.
@@
Recently, advancements in point-based differentiable rendering techniques (Kerbl et al., 2023; Held et al., 2024; Huang et al., 2024a, b) have demonstrated notable success in enhancing both rendering speed and high-fidelity reconstruction by leveraging smooth primitives within rasterization-based pipelines. While these point-based rendering techniques offer numerous advantages, they inherently represent scenes as unstructured point clouds. In contrast, contemporary graphics pipelines are predominantly designed for structured representations, such as polygonal meshes. These structured forms are widely adopted in a variety of applications, including animation (James and Twigg, 2005; Sorkine and Alexa, 2007), physical simulation (Ando et al., 2013; Baraff and Witkin, 2023; Liu et al., 2013), and editing (Yu et al., 2004; Zorin et al., 1997), where an explicit and coherent geometric structure is essential. Consequently, the absence of such inherent object structure in point-based methods significantly constrains their broader applicability.
@@
Our primary aim in this study is to bridge the gap between recent high-fidelity differentiable point-based representations and traditional graphics pipelines tailored for structured meshes. To this end, we introduce StructuredField, a novel method that unifies high-fidelity differentiable rendering and high-quality structured geometry. Our fundamental representation within StructuredField is the tetrahedral mesh, as it is suited for volumetric data and has been widely utilized in volume rendering (Ma and Crockett, 1997; Yagel et al., 1996; Weiler et al., 2003). To enable high-fidelity visual reconstruction, StructuredField reparameterizes the attributes of 3D Gaussian primitives with the geometric parameters of the underlying tetrahedral mesh. This strategy facilitates the direct, end-to-end differentiable optimization of the tetrahedral mesh using image-based losses, thereby achieving detailed reconstruction and high-quality rendering.
However, directly optimizing the vertex positions of tetrahedral meshes can readily lead to anomalous geometric structures, such as self-intersections or element inversions. To prevent these issues and ensure the integrity of the geometric structure throughout the optimization process, we analyze two main causes of mesh anomalies: self-intersections and element inversions, and propose a novel orientation-preserving homeomorphism (Dinh et al., 2017; Behrmann et al., 2019), which constrains the feasible deformation space of the mesh vertices, strictly maintaining a valid, intersection-free, and inversion-free tetrahedral mesh.
Furthermore, the capacity of a 3D representation to capture multi-scale details, like multi-resolution hash encodings in Instant-NGP (Müller et al., 2022) and adaptive density control in 3D Gaussian Splatting (Kerbl et al., 2023), is fundamental for achieving high-fidelity reconstruction. However, when extending to tetrahedral representations, achieving operations similar to adaptive density control often requires explicit, multi-level subdivision of the tetrahedra. Such explicit subdivision inherently introduces non-conformal interfaces between tetrahedral elements, which are problematic for downstream applications like physical simulation. Instead, we utilize an implicit subdivision strategy that adaptively refines regions for rendering by introducing implicit child tetrahedra without modifying the fixed topology of the initial conformal mesh. Consequently, by applying these two strategies, our representation inherently ensures high-quality geometry for simulation while providing sufficient expressive power for high-fidelity reconstruction.
@@
  • We introduce StructuredField, a novel 3D representation that seamlessly combines structured tetrahedral geometry with high-fidelity differentiable rendering capabilities through the reparameterization of 3D Gaussian functions.
@@
  • We design a novel orientation-preserving homeomorphism constraint to ensure that the tetrahedral mesh remains inversion-free and self-intersection-free during optimization.
  • We propose an implicit subdivision strategy for tetrahedral meshes that maintains a high-quality base mesh structure suitable for physical simulations while ensuring adaptive multi-scale details for high-fidelity reconstructions.
Figure 2. Overview of StructuredField. Given multi-view images as input, we reconstruct the 3D scene using a structured tetrahedral mesh. Our implicit multi-level subdivision module iteratively refines the initial input tetrahedra, creating a hierarchical structure. The root nodes of this hierarchy (the base tetrahedra) form a high-quality geometric foundation; their vertex positions are optimized using an orientation-preserving homeomorphism to maintain an inversion-free and robustly structured geometry, which supports subsequent applications. The leaf node tetrahedra, capturing the finest adaptive details, are then reparameterized into 3D Gaussian Splatting (3DGS) primitives for high-fidelity differentiable rendering. Based on our structured mesh representation, the reconstructed model can be further applied to physical simulations, deformation, and other graphics applications.
2. Related work
2.1. 3D Reconstruction
@@ 2. Related work
Reconstructing 3D scenes from multi-view images is a longstanding problem in both computer graphics and computer vision. Traditional 3D reconstruction techniques include Structure-from-Motion (SfM) pipelines (Agarwal et al., 2009; Frahm et al., 2010) to estimate camera poses and obtain sparse point clouds, followed by surface reconstruction through dense multi-view stereo (Furukawa et al., 2010; Jin et al., 2005; Furukawa and Ponce, 2010). These methods rely on hand-crafted features to acquire fine textures and geometry, and struggle to reconstruct view-dependent colors. Significant advancements have been achieved in NVS, particularly since the introduction of Neural Radiance Fields (NeRF) (Mildenhall et al., 2020a). The original NeRF represents the scene as an MLP, which maps positional encodings of spatial locations and directions to attributes including color and density, and utilizes a volume rendering process to achieve realistic rendering. Various works have enhanced the performance of NeRF (Müller et al., 2022; Barron et al., 2023, 2021) or extended them to large scenes (Barron et al., 2022; Tancik et al., 2022). More recently, 3DGS (Kerbl et al., 2023) optimizes anisotropic 3D Gaussian primitives, demonstrating real-time photorealistic reconstruction results. This method has been rapidly extended to multiple domains (Lin et al., 2024; Keetha et al., 2024; Tang et al., 2024; Xiang et al., 2024). Despite these successes, point-based representations are unstructured, limiting their further applications. In this paper, we demonstrate detailed reconstruction while maintaining a structured tetrahedral mesh, and showcase its subsequent applications.
2.2. Mesh-based Representation
@@
Explicit representations have served as a cornerstone within 3D modeling and computer graphics for decades (Woo et al., 1999). Conventional geometric representations like point clouds, voxels and polygonal meshes have been extensively revisited in the context of 3D deep learning. Polygonal meshes are particularly attractive due to their structured geometry and efficient rendering properties. Recently, differentiable rendering methods (Liu et al., 2019; Pidhorskyi et al., 2025) leveraging mesh representations have enabled the production of high-quality renderings. However, the optimization processes inherent to mesh representations are often hampered by rigid topological constraints, resulting in limited flexibility and reduced capacity to accurately depict realistic appearances. Our method integrates inversion-free mapping to effectively regulate and maintain structural integrity during the mesh optimization process while ensuring high quality reconstruction, addressing these limitations.
@@
On the other hand, NeRF (Mildenhall et al., 2020b) has attracted considerable attention due to their ability to deliver high-fidelity renderings. Substantial research has focused on integrating implicit methods with mesh structures to extend and refine NeRF (Wang et al., 2021). For instance, some approaches (Wang et al., 2023b; Wan et al., 2023; Turki et al., 2024) utilize meshes to constrain the sampling regions within an adaptively learned narrow band of two explicit meshes. On the other hand, Tetra-NeRF (Kulhanek and Sattler, 2023; Rosu and Behnke, 2023) employs tetrahedral meshes as feature grids to accelerate the training process. Additionally, leveraging the fast rendering capabilities of meshes, Baked SDF and its variants (Yariv et al., 2023; Reiser et al., 2024) have baked radiance fields into meshes to facilitate real-time rendering. Mesh representations have also been pivotal in the deformation of radiance fields. NeRFShop and Cage-NeRF (Peng et al., 2022; Jambon et al., 2023) use meshes as cages to drive the deformation of radiance fields.
@@
Recently, point-based representations (Kerbl et al., 2023; Huang et al., 2024a, b) have emerged, utilizing smooth primitives to represent radiance fields and employing rasterization techniques to accelerate rendering. In subsequent works, Mani-gs and GaMes (Gao et al., 2024) employ high-quality triangle meshes or triangle soups as proxies, binding primitives to the triangle mesh to drive the deformation of 3D Gaussians. Additionally, VR-GS (Jiang et al., 2024) and D3GA (Zielonka et al., 2025) adopt tetrahedral meshes as proxies to facilitate the deformation of objects. However, these methods directly bind the mesh to Gaussians without optimizing the positions of the mesh vertices. As a result, any imperfections in the initial mesh can significantly degrade the final reconstruction quality. In contrast, our method reparameterizes Gaussians with tetrahedral mesh vertices and optimizes these vertices, thereby allowing for high-quality reconstruction.
2.3. Inversion-free Mesh Optimization
@@
Maintaining mesh quality during the optimization process remains a challenging problem. Element inversion is a major factor causing mesh quality issues (Fu et al., 2021). Traditional methods (Schüller et al., 2013; Rabinovich et al., 2017) impose constraints on the Jacobian of the mapping during optimization to ensure the Jacobian remains positive, thereby enforcing that the mapping is orientation-preserving. In addition to being inversion-free, intersection-free boundaries are also necessary during the optimization. This is typically achieved by constraining the mapping to be bijective. Traditional methods (Jiang et al., 2017; Misztal and Bærentzen, 2012) introduce scaffold meshes to convert the globally overlap-free constraint into a locally flip-free condition. Recently, invertible networks (Behrmann et al., 2019; Dinh et al., 2017) offer another solution, leading to their widespread adoption in recent 3D deformation tasks. For instance, NDR (Cai et al., 2022) and Cadex (Lei and Daniilidis, 2022) employ invertible networks to model the motion of objects, thereby achieving dynamic object reconstruction. NFGP (Yang et al., 2021) utilizes invertible networks to perform geometric processing on implicit surfaces. Compared to these methods, we propose a novel invertible network that implicitly and strictly constrains the mapping to be inversion-free and intersection-free, and apply it during the optimization process to guarantee the mesh quality.
3. Method
@@
Given multi-view posed images of a scene, our primary objective is to reconstruct it with high fidelity and high-quality geometry. This enables seamless integration of the reconstructed 3D scene into existing computer graphics pipelines for applications such as rendering, animation, physical simulation, and deformation. To this end, we first introduce a novel tetrahedral mesh-based 3D representation that is structurally organized and facilitates differentiable optimization (Sec. 3.1). The mesh is optimized using an orientation-preserving homeomorphism to ensure high-quality structured geometry (Sec. 3.3). We further introduce a hierarchical implicit subdivision scheme to achieve conformal geometry while enabling multi-scale rendering detail (Sec.  3.2). An overview of our representation and reconstruction pipeline is shown in Fig. 2.
3.1. Representation of StructuredField
@@
Given a tetrahedral mesh $\mathcal{M}=(\mathcal{V},\mathcal{T})$, where $\mathcal{V}=\{v_{i}\}_{i=1}^{N}$ represents the set of vertices, $\mathcal{T}=\{T_{k}\}_{k=1}^{K}$ is the set of tetrahedron. Each tetrahedron $T_{k}\in\mathcal{T}$ is defined by four vertices $\{v_{k_{1}},v_{k_{2}},v_{k_{3}},v_{k_{4}}\}$. Tetrahedral meshes are generally challenging to render in a differentiable manner due to their discrete, non-smooth nature.
@@
To address this limitation, we build upon recent point-based differentiable rendering techniques, such as 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023), 2D Gaussian Splatting (2DGS) (Huang et al., 2024b), and convex splatting (Held et al., 2024). In our method, we establish a one-to-one correspondence between each tetrahedron and a primitive, using a differentiable reparameterization function $\mathbf{F}_{r}(\cdot)$ to map the parameters $T_{k}$ of the tetrahedron to the parameters $\Theta_{k}$ of the primitive:
(1) $\Theta_{k}=\mathbf{F}_{r}\bigl{(}T_{k}\bigr{)}.$
@@
The gradients are propagated back to the parameters of tetrahedra through the differentiable reparameterization function, enabling the optimization of vertex positions and associated attributes in a fully differentiable framework. In the following, we demonstrate how the parameters of 3DGS can be reparameterized using the parameters of a tetrahedral mesh.
Figure 3. Data Structure of Hierarchical Implicit Subdivision. Base tetrahedra can be iteratively subdivided in a 1-to-4 manner, forming a quadtree-like hierarchy. The root node tetrahedra constitute a high-quality and conformal geometry, designated for downstream applications such as physical simulation. Leaf node tetrahedra are utilized for high-quality rendering. The positions of the implicitly defined subdivision control points are optimized via their barycentric coordinates within their respective parent tetrahedra, while the positions of the root node vertices are optimized using our Orientation-preserving Homeomorphism.
@@
3DGS (Kerbl et al., 2023) represents a scene with a set of 3D Gaussians $\mathcal{G}=\{\boldsymbol{g}_{k}\}_{k=1}^{K}$, where each Gaussian $\boldsymbol{g}_{k}$ encodes the following attributes: mean $\boldsymbol{\mu}_{k}\in\mathbb{R}^{3}$, scales $\boldsymbol{s}_{k}\in\mathbb{R}^{3}$, rotation $\boldsymbol{r}_{k}\in\mathbb{R}^{4}$, color $\mathbf{c}_{k}\in\mathbb{R}^{3}$, and opacity $o_{k}\in\mathbb{R}$. In our representation, each tetrahedron $T_{k}$ contains four vertices with positions $\{\mathbf{v}_{k_{i}}\}_{i=1}^{4}$, spherical harmonic coefficients $\{\mathbf{SH}_{k_{i}}\}_{i=1}^{4}$, and weights $\{w_{k_{i}}\}_{i=1}^{4}$. We first compute the mean $\boldsymbol{\mu}_{k}$ and an initial, PCA-derived covariance matrix $\boldsymbol{\Sigma}_{k}^{\prime}$ from the weighted vertex positions of $T_{k}$, following a PCA-style approach (Abdi and Williams, 2010):
(2) $\boldsymbol{\mu}_{k}=\frac{\sum_{i=1}^{4}w_{k_{i}}\,\mathbf{v}_{k_{i}}}{\sum_{i=1}^{4}w_{k_{i}}},\quad\boldsymbol{\Sigma}_{k}^{\prime}=\sum_{i=1}^{4}w_{k_{i}}\bigl{(}\mathbf{v}_{k_{i}}-\boldsymbol{\mu}_{k}\bigr{)}\bigl{(}\mathbf{v}_{k_{i}}-\boldsymbol{\mu}_{k}\bigr{)}^{\top}.$
This $\boldsymbol{\Sigma}_{k}^{\prime}$ captures the initial anisotropic shape implied by the vertex distribution of $T_{k}$. The optimizable quaternion $\mathbf{q}_{k}$ is converted to its corresponding rotation matrix $\Delta\mathbf{R}_{k}$. The final anisotropic covariance matrix $\boldsymbol{\Sigma}_{k}$ for the 3D Gaussian is then obtained by applying this learned rotation:
(3) $\boldsymbol{\Sigma}_{k}=\Delta\mathbf{R}_{k}\boldsymbol{\Sigma}_{k}^{\prime}\Delta\mathbf{R}_{k}^{T}.$
This explicit, optimizable rotation allows the 3D Gaussian primitive to more flexibly orient itself to fit the local scene content associated with $T_{k}$.
The opacity attribute $o_{k}$ is defined on the tetrahedron instead of on the vertices, so the opacity of its corresponding Gaussian is equivalent to the opacity of the tetrahedron. The color of the Gaussian is obtained by a weighted average of the colors computed from the spherical harmonic function of each vertex:
(4) $\mathbf{c}=\frac{\sum_{i=1}^{4}w_{i}\mathbf{c}_{i}}{\sum_{i=1}^{4}w_{i}},\quad\mathbf{c}_{i}=\mathbf{SH}_{i}(\mathbf{d}),$
@@
After reparameterizing 3D Gaussians with tetrahedral mesh, the scene can be rendered in a differentiable manner. However, freely optimizing the tetrahedral vertices leads to low-quality tetrahedral mesh, as shown in Fig. 6. In Section 3.3, we introduce the orientation-preserving homeomorphism to address this issue.
3.2. Hierarchical Implicit subdivision